# Bootstrapping World Models from Dynamics Models in Multimodal Foundation Models

### v1

@@ 3.3 Test-time Verification
Finally, we introduce an inference-time strategy which harnesses the CDM as a verifier to enhance CWM performance. Inspired by recent work on scaling test-time compute (muennighoff2025s1, ; snell2024scaling, ), we let the CWM generate $N$ candidate observations. Each candidate is paired with the source and scored by the CDM, which assigns each a predicted likelihood, interpreted as a reward. The final prediction of the CWM is selected by maximising the CDM’s reward:
$\hat{o}_{t}=\mathop{\text{argmax}}_{i\in\{1,\dots,N\}}p_{\text{CDM}}\left(a\mid o_{s},o_{t}^{(i)}\right),\quad\text{where }o_{t}^{(i)}\sim p_{\text{CWM}}(o_{t}\mid o_{s},a),$
where $\hat{o}_{t}$ is the selected prediction.
@@ 4.1 Experimental Setting
We select Aurora-Bench (krojer2024aurora, ) for evaluation of both dynamics and world models. This dataset provides high-quality data for action-centric edits, covering a wide array of phenomena and assessing a model’s alignment with the physical world, including temporal and spatial reasoning. We choose 5 subsets: MagicBrush for specialised image editing, Action-Genome (AG) and Something-Something (Something) for real-world actions and scenarios. Whatsup focuses on spatial reasoning, whereas Kubric contains synthetic samples from a physical engine (greff2022kubric, ).
Baselines.
We report Chameleon’s zero-shot performance (C-ZS). We also fine-tune Chameleon on Aurora’s training set as our first baseline (C-FT). We compare CWM with C-FT in both a single-prediction setting and in a best-of-$N$ setting. The latter provides a ceiling performance for inference-time verification with CDM. Additionally, we include three state-of-the-art diffusion models specialised for image editing as baselines, such as PixInstruct (brooks2023instructpix2pix, ), GoT (fang2025got, ) and SmartEdit (huang2024smartedit, ). As a sanity check, we also report the metric scores obtained by simply copying the source observation input as the next-observation prediction (Copy).
@@
For next-observation prediction evaluation, following (fang2025got, ), we rely on GPT4o-as-a-judge as it is the only metric that reliably penalises Copy. In Appendix A.3, we show four other metrics, e.g., CLIP, which assign high scores to Copy. GPT4o-as-a-judge scores consider two criteria, one for the editing success rate and one for visual consistency with the original. We take the minimum of the two as the final score. The prompt for GPT4o-as-a-judge is provided in Appendix A.6.
4.2 Chameleon Dynamics Model
@@
Table 1: Performance of dynamics models performance on action prediction, measured by text similarity metrics: BERTScore (BS; (zhangbertscore, )), ROUGE-1/2/L (R-1, R-2, R-L; (lin2004rouge, )) and BLEU (papineni2002bleu, ).
BS R-1 R-2 R-L BLEU
VILA-U Fine-tuned 0.40 0.38 0.20 0.37 0.15
Chameleon Zero-Shot (C-ZS) 0.05 0.09 0.02 0.08 0.00
Chameleon Fine-Tuned (CDM) 0.40 0.39 0.20 0.37 0.17
Chameleon Fine-Tuned (CDM) + DS 0.45 0.45 0.27 0.44 0.20
Figure 4: Comparison of negative log-likelihoods (lower values indicate stronger model preference) of the action predicted by CDM for ground-truth trajectories versus four types of negative trajectories.
We evaluate the dynamics models based on the textual similarity of the predicted action with the ground-truth action in Aurora-Bench, as shown in Table 1. Our results demonstrate that fine-tuning is necessary to elicit Chameleon’s ability to verbalise the dynamics from two observations. We then compare Chameleon fine-tuned on action prediction (CDM) with the fine-tuned version of another state-of-the-art VLM, VILA-U (wu2024vilaU, ). CDM is on par or superior to VILA-U fine-tuned, justifying our choice of Chameleon as a foundation model for our experiments. Table 1 also provides an ablation showing that downsampling trajectories from Kubric (DS) in the training data further boosts CDM performance (CDM + DS). This suggests that data sourced from simulations do not necessarily translate into better dynamics modelling in real-world examples. We use the DS version of CDM in the rest of the experiments as the best-performing dynamics model. In Figure 4, we further evaluate CDM on discriminating between ground-truth and negative trajectories, as in Section 2. Now, we observe that CDM is mostly successful in identifying manipulated trajectories as such, except for Copy. These results corroborate the feasibility of annotating actions for key-frame pairs.
4.3 Chameleon World Model
Table 2: Model performance on MagicBrush, AG, Something, WhatsUp, and Kubric from Aurora-Bench in terms of GPT-4o scores. For C-FT and CWM, we report their performance for both single prediction and best-of-N. The average scores for each model are shown at the bottom. We bold the best model overall for each subset and highlight the best and worst scores among our variants for each setting. SE: SmartEdit. Detailed results with 5 metrics are presented in Table 3.
Datasets Models
Copy PixInstruct GoT SE C-ZS C-FT +Best-of-3 CWM +Best-of-3
MagicBrush 0.000 3.120 5.960 6.710 0.000 2.520 3.270 3.920 3.920
AG 0.000 1.200 1.610 3.080 0.170 2.480 2.740 3.640 3.640
Something 0.000 0.957 2.620 2.810 0.370 3.110 3.110 2.920 3.310
WhatsUp 0.000 0.000 1.580 0.755 0.146 0.880 0.980 0.540 0.540
Kubric 0.000 1.880 3.920 3.700 0.140 7.300 7.300 7.320 7.780
Average 0.000 1.430 3.140 3.410 0.165 3.260 3.480 3.670 3.840
Figure 5: Detailed scores of GPT4o-as-a-judge evaluation for loss-weighting and standard training. We report the scores for Editing Success (ES) and Minimal Editing (ME). MB: MagicBrush, AG: Action-Genome, ST: Something-Something, WU: WhatsUp, KU: Kubric. We highlight the best and worst scores for each category.
@@ 4.3 Chameleon World Model
MB 3.73 8.17 3.68 8.46
AG 3.18 8.03 2.37 8.13
ST 3.32 7.01 2.78 7.20
WU 0.54 7.25 0.76 7.19
KU 7.75 8.49 7.24 8.70
Avg. 3.71 7.80 3.37 7.94
GPT4o 3.67 3.58
Automatic evaluation.
Next, we test CWM on next-observation prediction for each of the Aurora-Bench subsets, reporting GPT4-as-a-judge scores in Figure 2. We first notice that the state-of-the-art image editing models (i.e., PixInstruct, GoT, SmartEdit) tend to specialise in the image editing benchmark, MagicBrush ($5.96$ and $6.71$ GPT4o scores for GoT and SmartEdit). Nevertheless, in the action-centric subsets, including Action-Genome (AG), Something and Kubric, they are mostly behind CWM and even C-FT. In particular, CWM outperforms all other models in these 3 subsets, achieving gains of 18%, 4%, and 86%, respectively, over the best diffusion baselines. In addition, it boosts the highest average performance across subsets, with an 8% increase. Crucially, comparing CWM and C-FT reveals the benefit of augmenting the training data with synthetic triplets bootstrapped from the CDM, as it yields a 13% performance margin. CWM also outperforms C-FT on the best-of-N setting (amini2024best-of-n, ), indicating the potential for inference-time verification as best-of-N is effectively an oracle for its performance.
@@
Following (krojer2024aurora, ), we conduct a blind human evaluation comparing GoT, SmartEdit, C-FT, and our proposed CWM. We randomly sample 5 examples from each subset within Aurora-Bench and present the outputs generated by each of the four models. Human annotators are asked to identify the best and worst generated observations based on three criteria: (1) Realism: the generated image should exhibit natural textures and lighting while remaining faithful to the input scene; (2) Instruction-Following Ability: the edit should clearly reflect the given instruction; and (3) Over-Editing: the modification should be minimal and focused, altering only what is necessary. Each model receives +1 point for being selected as the best, -1 for the worst, and 0 otherwise. We compute the average scores over 350 annotated samples, as reported in Table 7. The results align with automatic evaluations: image-editing models excel in the MagicBrush domain, but fall short on action-centric datasets such as Action-Genome, Something-Something, and Kubric. In contrast, CWM outperforms C-FT on all three of these datasets, highlighting its strength in next-observation prediction in real-world, action-centric trajectories.
Figure 6: Ablation study of synthetic trajectories (Synth.) and loss weighting (LW) in CWM. Numbers are GPT-4o-as-judge scores ($\uparrow$, average of 3 runs). MB: MagicBrush, AG: Action-Genome, ST: Something-Something, WU: WhatsUp, KU: Kubric.
CWM w/o Synth. w/o LW
@@
[…]
GoT SE C-FT CWM
@@
To assess the importance of extra supervision from CDM-synthetic trajectories, Table 7 reports GPT-4o’s scores for this ablation. We see performance drops on most datasets—particularly on Something and AG—when the additional training data from unlabelled videos is removed, highlighting the effectiveness of bootstrapping CWM with large-scale real-world data via CDM. An exception is the WhatsUp dataset, which focuses on specific actions within a fixed scene; in this case, training in an open-domain setting may not transfer effectively.
@@
Based on Table 7, we also observe consistent degradation when loss weighting is removed, demonstrating the benefit of explicitly incorporating the recognition model into visual next-token prediction. To better understand the effect of loss weighting, Table 5 reports the average scores for two criteria used in the GPT-4o-as-a-judge evaluation separately: Editing Success (ES), which measures how well the model captures the intended action and performs the corresponding edit, and Minimal Editing (ME), which assesses whether the model introduces unnecessary modifications. The full distribution of GPT-4o scores is provided in Appendix A.7. Our analysis reveals that the primary bottleneck for CWM remains its ability to reliably follow the instruction, as reflected by the fact that ES scores are significantly lower than ME scores. Loss weighting partly solves this problem, increasing the editing success and reducing copying behaviour, albeit at the cost of sometimes over-editing the source observation.
@@
Figure 8: GPT-4o scores for test-time verification with $K$ samples, where $K\in\{1,2,4,8\}$. We use a blue line for C-FT and a red line for CWM, plotting the standard deviation as the shaded area. We indicate the scores for GoT (GT) and SmartEdit (SE) as horizontal lines.
Verification at Test Time.
We evaluate test-time verification using CDM in Figure 8, comparing C-FT and CWM with $K\in{1,2,4,8}$. Each experiment is repeated three times, and we visualise the mean and standard deviation. By increasing exploration on more candidate next observations, C-FT benefits from test-time verification on most datasets with real-world trajectories (e.g., AG, Something, WhatsUp), suggesting the effectiveness of CDM’s trajectory preferences. Increasing $K$ does not always improve performance (MagicBrush, Kubric), suggesting that bootstrapping with a dynamics model that shares the same foundation model backbone may be limiting. In contrast, CWM shows no clear gain, likely because it was trained with the synthetic trajectories and has already internalised CDM’s preferences—as is evident from its strong $K=1$ performance. In summary, CDM-based verification boosts C-FT’s performance to the same level as CWM, by leveraging more diverse samples at inference time rather than during training.
Qualitative Example. Figure 9 presents a real-world example demonstrating that CWM’s predicted observations can be guided through language expressing actions. CWM is also capable of iteratively generating future observations in multiple steps while maintaining consistency with previous frames.
@@
Figure 9: A qualitative case of real-world observation prediction, demonstrating CWM’s ability to steer predictions using language and perform sequential predictions. More cases from Aurora-Bench are in Appendix A.4.
@@ 5 Related Work
Despite the surge in interest for world modelling (2018_ha_schmidhuber_recurrent_world_model, ; sutton1988worldmodeling, ; hafner2019learningLatentDynamics, ), previous works focused mostly on building specialised ad-hoc world models. These world models can be explicitly learnt as a visual simulator (agarwal2025cosmos, ; bruce2024genie, ; brooks2024videoGenerationWorldSimulator, ), or enable planning with model-based reinforcement learning (hafner2019learning, ; micheli2022transformers, ; robine2023transformerworldmodelTWM, ; alonso2024diffusionForWorldModeling, ; Hafner2025, ). Instead, we focus on leveraging large-scale multimodal foundation models (lin2024video, ; chen2025janus, ; wu2024vilaU, ) to develop world models, which is more attractive due to the inductive bias they provide from their extensive training. This is possible thanks to frameworks that integrate observations, actions, and rewards into a unified sequence of tokens in autoregressive Transformers (NEURIPS2024_7dbb5bfa, ), building on pioneering works such as Decision Transformers (chen2021decisionTransformers, ) and GATo (reed2022generalistGATo, ). Related to our work, (chen-vlm-representation-for-RL, ) initialise the parameters of RL policies with VLMs, thus taking advantage of the abundant and general world knowledge encoded in their representations. 3D-VLA (zhen20243dWorldModelVLM, ) introduces a set of interaction tokens into a Large Language Model to engage with the environment as an embodied agent. (yang2024video, ; soni2024videoagent, ) explore large-scale self-supervised learning via next token or frame prediction to build a unified model absorbing internet knowledge, learning from interaction via video.
@@
Aurora-Bench (krojer2024aurora, ) was the first to approach world modelling through the lens of an action-centric image editing task. With advanced native VLMs capable of the interleaved generation (team2024chameleon, ; chern2024anole, ), we systematically investigate how this data may help us bootstrap a world model, implicitly stored in the VLMs, with an easier-to-train dynamics model. Most similar to our work, (baker2022videoPretraining-vpt, ) train a dynamics model which aims to uncover the underlying action between video frames in unlabelled video frames from the Minecraft game. Through this model, they synthesise trajectories to train a policy for sequential decision making. In contrast with (baker2022videoPretraining-vpt, ), we focus on next-observation prediction as a task to evaluate world modelling. First, this allows us to port the observation space to real-world frames, rather than simulated ones, hence assessing whether world models are well aligned with the physical environment. Second, this broadens the space of actions from a few choices to the combinatorially infinite and expressive space of language, capturing a significantly more diverse range of dynamics.
6 Conclusion
In this work, we explored whether we can develop world models from VLMs. By evaluating them on action-centric image editing Aurora-Bench (krojer2024aurora, ), we first show that these models lack a clear preference for ground-truth real-world trajectories. To address this, we induce a dynamics model from the same VLM to bootstrap a world model using automatic annotation of unlabelled real-world videos and inference-time verification. Experiments confirm the effectiveness of both strategies, with our general-purpose world model achieving state-of-the-art performance compared to existing approaches specialised for image editing.
Appendix A Appendix
A.1 Limitations

### v2

@@ 3.3 Test-time Verification
Finally, we introduce an inference-time strategy which harnesses the IDM as a verifier to enhance FDM performance. Inspired by recent work on scaling test-time compute (Muennighoff et al., 2025; Snell et al., 2024), we let the FDM generate $N$ candidate observations. Each candidate is paired with the source and scored by the IDM, which assigns each a predicted likelihood, interpreted as a reward. The final prediction of the FDM is selected by maximising the IDM’s reward:
$\hat{o}_{t}$ $=\mathop{\mathrm{argmax}}_{i\in\{1,\dots,N\}}p_{\text{IDM}}\!\left(a\mid o_{s},o_{t}^{(i)}\right),$
$\quad\text{where }o_{t}^{(i)}\sim p_{\text{FDM}}(o_{t}\mid o_{s},a).$
where $\hat{o}_{t}$ is the selected prediction. While this strategy is training-free, it requires sampling multiple candidates at inference time.
@@ 4 Experiments and Results
In Section 4.1, we first describe the experimental setup, including benchmarks, baselines, and evaluation metrics. We then report results on the inverse dynamics prediction task in Section 4.2 to verify that fine-tuning on a limited amount of examples is sufficient to develop robust IDP in a VLM. This is followed in Section 4.3 by automatic and human evaluation of both strategies introduced in Section 3 to bootstrap FDP, as well as ablation studies. Finally, we analyse inference-time verification and examine the transfer of forward dynamics prediction to two spatial reasoning benchmark.
@@ 4.1 Experimental Setting
We select Aurora-Bench (Krojer et al., 2024) for evaluation of both IDM and FDM. This dataset provides high-quality data for action-centric edits, covering a wide array of phenomena and assessing a model’s alignment with the physical world. We choose 5 subsets: MagicBrush for specialised image editing, Action-Genome (AG) and Something-Something (Something) for real-world actions. Whatsup focuses on spatial reasoning, whereas Kubric contains samples from a physical engine (Greff et al., 2022).
Models and Baselines.
We experiment with two unified VLMs, Chameleon-7B (Chameleon Team, 2024) and Liquid-8B (Wu et al., 2024b). The variants of these models fine-tuned on both supervised and weakly supervised data with loss weighting (i.e., with the training-time bootstrapping strategy) are indicated as C-FDM and L-FDM, respectively. As our first baselines, we use the same two VLMs either zero-shot (C-ZS and L-ZS) or fine-tuned only on Aurora’s supervised data (C-FT and L-FT). Additionally, we include three state-of-the-art diffusion models specialised for image editing as baselines, such as PixInstruct (Brooks et al., 2023), GoT (Fang et al., 2025) and SmartEdit (Huang et al., 2024). Finally, as a sanity check, we also report the metric scores obtained by simply copying the source observation input as the next-observation prediction (Copy).
@@
For next-observation prediction (FDP) evaluation, following Fang et al. (2025), we rely on GPT4o-as-a-judge as it is the only metric that reliably penalises Copy. In Appendix B, we show four other metrics, e.g., CLIP, which assign high scores to Copy. GPT4o-as-a-judge scores consider two criteria, one for the editing success rate and one for visual consistency with the original. We take the minimum of the two as the final score. The prompt for the judge is provided in Appendix F.
4.2 Inverse Dynamics Prediction
BS R-1 R-L BLEU
Chameleon ZS 0.05 0.09 0.08 0.00
Chameleon IDM 0.45 0.45 0.44 0.20
Liquid IDM 0.41 0.41 0.36 0.21
@@ 4.2 Inverse Dynamics Prediction
Table 1: Performance of Inverse Dynamics Models on action prediction, measured by text similarity metrics: BERTScore (BS; Zhang et al. 2020), ROUGE (R-1, R-L; Lin 2004) and BLEU (Papineni et al., 2002).
We evaluate the IDM based on the textual similarity of the predicted action with the ground-truth action in Aurora-Bench. Results are shown in Table 1. Our results demonstrate that a moderate amount of fine-tuning is necessary to let unified VLMs verbalise the dynamics connecting two observations, as evidenced by the large It is also worth noting that both models’ IDP performance (in terms of BLEU) is similar. We leverage these IDP versions of Chameleon and Liquid to bootstrap FDP, whose results are reported in Section 4.3.
4.3 Forward Dynamics Prediction
Next, we report FDP results for each of the two bootstrapping strategies: in Section 4.3.1 for trajectory synthesis and in Section 4.3.2 for inference-time verification.
4.3.1 Synthesising Trajectories with IDP
Dataset Models
PixInst GoT SE C-ZS C-FT C-FDM L-ZS L-FT L-FDM
Forward-dynamics Prediction (GPT4o score)
[…]
Spatial Reasoning (Accuracy)
SpatialMQA 26.1 25.8 27.2 27.7 28.0 27.8
EmbodiedSpatial-Bench 15.1 21.2 17.5 32.6 33.2 33.8
Table 2: Performance on Aurora-Bench (GPT-4o-as-a-judge score) and spatial reasoning benchmarks (accuracy). We bold the best model per dataset. Among our variants, we highlight the best and worst scores. SE: SmartEdit. Greedy decoding is used for spatial reasoning evaluation and applied only to VLMs.
Automatic evaluation.
To test trajectory synthesis, we evaluate FDM on next-observation prediction for each of the Aurora-Bench subsets, reporting GPT4o-as-a-judge scores in Table 2. We first notice that the state-of-the-art image editing models (i.e., PixInstruct, GoT, SmartEdit) tend to specialise on the proper image editing subset MagicBrush ($5.96$ and $6.71$ GPT4o scores for GoT and SmartEdit). Nevertheless, in the action-centric subsets, including Action-Genome (AG), Something and Kubric, they are mostly behind bootstrapped models (C/L-FDM) and even the fine-tuned baselines (C/L-FT). Crucially, considering the average performance of C/L-FDM on all subsets of Aurora-Bench reveals the benefit of augmenting the training data with synthetic triplets bootstrapped from the IDM (vs. FT), as it yields a 13% gain for Chameleon and 7% for Liquid, and the benefit of fine-tuning off-the-shelf VLMs for FDP more broadly (vs. ZS).
@@ 4.3.1 Synthesising Trajectories with IDP
[…]
Table 3: Detailed scores of GPT4o-as-a-judge evaluation for loss-weighting and standard training. We report the scores for Editing Success (ES) and Minimal Editing (ME). MB: MagicBrush, AG: Action-Genome, ST: Something-Something, WU: WhatsUp, KU: Kubric. We highlight the best and worst scores for each category. We report the average of ES and ME as Avg. and the final score as GPT4o.
@@
Following Krojer et al. (2024), we conduct a blind human evaluation comparing GoT, SmartEdit, C-FT, and C-FDM. We randomly sample 5 examples from each subset within Aurora-Bench and present the outputs generated by each of the four models. Human annotators are asked to identify the best and worst generated observations based on three criteria: (1) Realism: the generated image should exhibit natural textures and lighting while remaining faithful to the input scene; (2) Instruction-Following Ability: the edit should clearly reflect the given action; and (3) Over-Editing: the modification should be minimal and focused, altering only what is necessary to fulfil the action. Each model receives +1 point for being selected as the best, -1 for the worst, and 0 otherwise. We compute the average scores over 350 annotated samples, as reported in Table 5. The results are well aligned with automatic evaluations in Table 2: image-editing models excel in the MagicBrush and WhatsUp subsets, but fall short on action-centric datasets such as Action-Genome, Something-Something, and Kubric. In contrast, C-FDM outperforms C-FT (and all other baselines) on all three of these datasets, highlighting its strength in next-observation prediction in real-world, action-centric trajectories despite remaining a general-purpose VLM.
C-FDM w/o Synth. w/o LW
@@
Table 4: Ablation study of synthetic trajectories (Synth.) and loss weighting (LW) in C-FDM. Numbers are GPT-4o-as-judge scores ($\uparrow$, average of 3 runs). MB: MagicBrush, AG: Action-Genome, ST: Something-Something, WU: WhatsUp, KU: Kubric.
GoT SE C-FT C-FDM
@@
Table 5: Human evaluation results. $\dagger$} indicates all results whose gap with respect to C-FDM is significant, based on a Wilcoxon signed-rank test ($p=0.05$). MB: MagicBrush, AG: Action-Genome, ST: Something-Something, WU: WhatsUp, KU: Kubric.
@@
To assess the importance of extra supervision from IDM-synthetic trajectories, Table 5 reports GPT-4o’s scores for this ablation. We see performance drops on most datasets—particularly on Something and AG—when the additional training data from unlabelled videos is removed, highlighting the effectiveness of bootstrapping C-FDM with large-scale real-world data via IDM. An exception is the WhatsUp dataset, which focuses on specific actions within a fixed scene; in this case, training in an open-domain setting may not transfer effectively.
@@
Based on Table 5, we also observe consistent degradation when loss weighting is removed, demonstrating the benefit of explicitly incorporating the recognition model into visual next-token prediction. To better understand the effect of loss weighting, Table 3 reports the average scores for two criteria used in the GPT-4o-as-a-judge evaluation separately: Editing Success (ES), which measures how well the model captures the intended action and performs the corresponding edit, and Minimal Editing (ME), which assesses whether the model introduces unnecessary modifications. The full distribution of GPT-4o scores is provided in Appendix G. Our analysis reveals that the primary bottleneck for C-FDM remains its ability to reliably follow the instruction, as reflected by the fact that ES scores are significantly lower than ME scores. Loss weighting partly solves this problem, increasing the editing success and reducing copying behaviour, albeit at the cost of sometimes over-editing the source observation.
@@
Figure 4: GPT-4o scores for test-time verification with $K$ samples, where $K\in\{1,2,4,8\}$. We use a blue line for C-FT and a red line for L-FT. For C-FT, we plot the standard deviation as the shaded area due to its large variance. We indicate the scores for GoT (GT) and SmartEdit (SE) as horizontal lines.
Image Editing as an Auxiliary Task.
Finally, we assess whether enhanced FDP capabilities are beneficial for a broader range of vision–language tasks. Since Aurora and the action-annotated videos contain diverse spatial relations (e.g., left/right orientation), we evaluated whether this supervision helps VLMs generalise beyond editing. We tested our models on two spatial reasoning benchmarks: SpatialMQA (Liu et al., 2025) and EmbodiedSpatial-Bench (Du et al., 2024), and report the corresponding accuracy in Table 2. Both Chameleon and Liquid trained with the FDP objective outperform the zero-shot baseline, demonstrating that the FDP task transfers beyond action-centric editing and highlighting FDP as a signal for enhancing spatial reasoning.
@@
Figure 5: A qualitative case of real-world next-observation prediction, demonstrating C-FDM’s ability to steer predictions using language and perform sequential predictions. More cases from Aurora-Bench are in Appendix C.
4.3.2 Inference-time Verification
We evaluate our test-time strategy using IDM to choose among candidates generated by C/L-FT in Figure 4, using $K\in{1,2,4,8}$. By increasing exploration on more candidate next observations, Chameleon and Liquid benefit from test-time verification on most datasets with real-world trajectories (e.g., AG, Something, WhatsUp), indicating the reliability of IDM’s trajectory preferences. Increasing $K$ does not always improve performance (MagicBrush, Kubric), however, suggesting that bootstrapping with IDM that shares the same foundation model backbone may be limiting. In summary, IDM-based verification boosts performance to a similar level as FDM, by leveraging more samples during inference rather than training.
Qualitative Example.
Figure 5 presents a real-world example demonstrating that C-FDM is also capable of iteratively generating future observations in multiple steps while maintaining consistency with previous frames.
@@ 5 Related Work
Despite the surge in interest for world modelling (Ha and Schmidhuber, 2018; Sutton, 1988; Hafner et al., 2019a), previous works focused mostly on building specialised ad-hoc world models. These world models can be explicitly learnt as a visual simulator (Agarwal et al., 2025; Bruce et al., 2024; Brooks et al., 2024), or enable planning with model-based reinforcement learning (Hafner et al., 2019b; Micheli et al., 2022; Robine et al., 2023; Alonso et al., 2024; Hafner et al., 2025). Instead, we focus on leveraging general-purpose, pre-trained vision-language models (Lin et al., 2024; Chen et al., 2025; Wu et al., 2024c) to develop world models, which is more attractive due to the inductive bias they provide from their extensive training.
This is possible thanks to frameworks that integrate observations, actions, and rewards into a unified sequence of tokens in autoregressive Transformers (Wu et al., 2024a), building on pioneering works such as Decision Transformers (Chen et al., 2021) and GATo (Reed et al., 2022b). Related to our work, Chen et al. (2024) initialise the parameters of RL policies with VLMs, thus taking advantage of the abundant and general world knowledge encoded in their representations. 3D-VLA (Zhen et al., 2024) integrates a set of interaction tokens into a Large Language Model to engage with the environment as an embodied agent. Yang et al. (2024) and Soni et al. (2024) explore large-scale self-supervised learning via next token or frame prediction to build a unified model absorbing internet knowledge, learning from interaction via video.
@@
Most similar to our work, Baker et al. (2022) train a dynamics model which aims to uncover the underlying action between video frames in unlabelled video frames from the Minecraft game. Through this model, they synthesise trajectories to train a policy for sequential decision making. In contrast with Baker et al. (2022), we focus on actin-centric next-observation prediction as a task to evaluate FDP. First, this allows us to port the observation space to real-world frames, rather than simulated ones, hence assessing whether VLMs can eventually be developed into world models. Second, this broadens the space of actions from a few choices to the combinatorially infinite and expressive space of language, capturing a significantly more diverse range of dynamics.
6 Conclusions
In this work, we explored whether unified vision–language models (VLMs) can be endowed with the ability to predict forward dynamics, i.e., the next observation in the environment (e.g., an image frame) given the past observation and an action (e.g., a textual instruction). We first show that these models lack a clear preference for ground-truth real-world action–observation trajectories compared with adversarially manipulated ones.
To address this, we leverage an inverse-dynamics model (IDM) fine-tuned from the same VLM, which consists instead of predicting actions taking place between observations and is easier to learn, to bootstrap a better forward-dynamics model (FDM). Specifically, the IDM can be used to 1) automatically annotate pairs of frames from unlabelled videos, which are then used for weakly supervised training of the FDM; or 2) verify the best sample among multiple candidates generated from the FDM at inference time. Experiments confirm the effectiveness of both strategies, with our general-purpose forward-dynamics model achieving state-of-the-art performance compared to existing approaches specialised for image editing.
Limitations